# Semantic-ICP: Iterative Closest Point for Non-rigid Multi-Organ Point Cloud Registration

### v2

@@ 3.5 Ablation Studies
W/o semantic matching + RegReg $2.42^{*}$ $90^{*} 0.0$0250.0025
Semantic w/o RegReg $\mathbf{1.15}^{*}$ $\mathbf{0.66}^{*}$ $0.12^{*}$
Semantic + $Reg_{grad}$ $\mathbf{1.15}^{*}$ $\mathbf{0.66}^{*}$ $0.0045^{*}$
Semantic + $Reg_{mag}$ $\mathbf{1.15}^{*}$ $\mathbf{0.66}^{*}$ $0.0089^{*}$
Semantic + $Reg_{els}$ $\underline{1.17}$ $\underline{0.67}$ $\underline{0.0025}$
Semantic + $Reg_{grad}+Reg_{mag}$ $\mathbf{1.15}^{*}$ $\mathbf{0.66}^{*}$ $0.0045^{*}$
Semantic + $Reg_{grad}+Reg_{els}$ $\underline{1.17}$ $\underline{0.67}$ $\mathbf{0.0024}$
Semantic + $Reg_{mag}+Reg_{els}$ $\mathbf{1.15}$ $\underline{0.67}$ $\underline{0.0025}$
Semantic + $Reg$ (Ours) $\underline{1.17}$ $\underline{0.67}$ $\underline{0.0025}$
@@ 3.6 Sensitivity on Elasticity Values
To evaluate the sensitivity of our method to the Young’s modulus $E$ and Poission’s ratio $\nu$, we perform SemICP on the AbdominalMRCT dataset with varying elastic values. Based on finite element simulation [46], typical $E$ for soft tissues and abdominal organs ranges from 1kPa to 100 kPa, except only the stomach that reaches 500kPa, and $\nu$ typically ranges from 0.400 to 0.499. In $E$ sensitivity analysis, we fixed $\nu=0.499$ and varied $E$ across {1kPa, 10kPa, 100kPa, 500kPa, 1MPa}. To also understand the influence of $\nu$, we fixed $E$ at 1kPa and varied $\nu$ in {0.350, 0.400, 0.450, 0.499}.
The sensitivity of SemICP to $E$ is presented in Table 8. As Young’s modulus $E$ increases, the surface matching errors HD95 and MSD increase slightly, while the spatial distortion SDLogJ decreases. However, the absolute differences across metrics remain small, particularly between 1kPa and 100kPa. The results indicate that as the tissue stiffness increases, the estimated deformation becomes more regularized, but its ability to accurately fit the boundary conditions slightly diminishes. However, in the soft tissues and soft organs range, the differences in performance are small. The sensitivity of SemICP to Poisson’s Ratio $\nu$ is shown in Table 9. The surface matching has a slight improvement when the $\nu$ decreases, while the spatial distortion SDLogJ increases. It indicates that assuming greater tissue compressibility weakens the linear energy regularization, allowing larger but less regularized deformations to fit the boundary conditions. Nonetheless, the differences in metrics remain small.
@@
HD95 $\downarrow$ 1.17 1.20* 1.23* 1.49* 1.85*
MSD $\downarrow$ 0.67 0.68* 0.71* 0.81* 0.91*
SDLogJ $\downarrow$ 0.0025 0.0018* 0.0011* 0.0008* 0.0008*
@@ 3.7 Sensitivity on Misalignment and Visualizable Ratio
To assess the sensitivity of our method and its robustness with respect to initial misalignments, we simulated partial observations and applied different rigid misalignments to the AbdominalMRCT dataset. We cropped the original point clouds to partially visible point clouds with 5%, 10%, 25% and 50% of points visible, and introduced additional Gaussian noise. We also added different levels of misalignment, with Euler rotation angles within $\pm 5^{\circ}$, $\pm 15^{\circ}$, and $\pm 30^{\circ}$. For each experiment, we generated 10 different simulated partial point clouds for each patient, resulting in a total of 80 synthetic point cloud pairs. We compare our SemICP-rigid method with ICP and GO-ICP. FR-ICP is excluded from comparison due to its higher HD95 in Table 3, indicating less favorable performance. As shown in Table 10, SemICP-rigid achieves lower HD95, MSD, and TRE compared to ICP and GO-ICP under challenging visibility conditions (5%, 10%, and 25%). When evaluating sensitivity to initial misalignment, both ICP and GO-ICP exhibit significantly increased registration error at rotation angles of $\pm 30^{\circ}$, indicating sensitivity to initialization. In contrast, SemICP-rigid demonstrates improved robustness, with smaller error increases under the same conditions. These results highlight that incorporating semantic labels enhances the accuracy and stability of the rigid registration step, even under limited visibility and challenging initial alignment.
@@
Figure 5: (a) Comparison of SemICP with and without semantic matching. Top: Without semantic+RegReg; Bottom: Semantic+RegReg. (b) Comparison of different regularization terms. Yellow indicates larger deformation; green and orange masks represent the liver and spleen in CT segmentation.
@@ 4 Discussion
Our work introduces a novel biomechanical energy-based regularization into point-matching-based point cloud registration. Existing methods (NR-ICP, AMM-NRR and Fast-RNRR) regularize the deformation with a local affine regularization to constrain deformation, which fails to capture the underlying biomechanical properties of medical point clouds. Moreover, these methods focus solely on surface matching, neglecting the fact that medical point clouds often undergo deformation as cohesive volumetric structures. This limits their ability to assess and preserve the underlying deformation topology. Our elastic energy-based regularization is physically motivated by the biomechanical properties of anatomical structures and may explain why it works better. Ablation studies confirm that the inclusion of the proposed elastic energy regularization leads to significant improvements in the plausibility and quality of the estimated deformation fields, as evidenced by reduced SDLogJ, with little degradation in the surface matching accuracy.
We do not apply the actual elasticity values in the regularization, but the sensitivity studies show that the differences in HD95, MSD and SDLogJ remain small with various $E$ and $\nu$. The non-rigid cost function includes two parts: the first part is boundary conditions that drive the alignment of the organ surfaces, and the second part is regularization that includes elastic energy. They work together to reach the equilibrium between accuracy and smoothness. Our sensitivity experiments on $E$ show that the registration is more likely to be driven by boundary conditions than by the regularization, particularly when $E$ is within the range of 1kPa to 100kPa, including the soft organs of interest like liver and kidney [47] and brain [48]. Similar results are observed in finite element-based registration. Chi et al. [49] found that a 30% material uncertainty results in registration errors within 1.3mm for thin and hollow structures, and though the maximum error for prostate can reach up to 4.5mm, most of the subvolumes have a much smaller registration error. Our sensitivity experiments on $\nu$ indicate that $\nu$ has a small influence on the registration results, which is similar to the finding in the finite element-based registration [46] that varying $\nu$ only leads to small changes in accuracy. In the future, inhomogeneous $E$ and $\nu$ can be employed based on the labels or information from elastography [33]. More complicated non-linear effects might be integrated, but it requires the consideration of the trade-off between algorithm complexity and accuracy. A research in finite element-based lung registration shows that the registration error is not significantly affected by the element type (elastic and hyperelastic) or linearity of analysis, suggesting the use of linear finite element analysis for simplicity [50].
@@
Our work requires semantic labels in the point cloud, and this paper focuses on registration without discussing the point cloud labeling process. Preoperative images are often labeled manually, but intraoperative labeling can hinder real-time applications. In future work, we plan to investigate deep-learning models to perform intraoperative segmentation for the full registration pipeline. Additionally, this method is designed for multi-organ registration, but it has the potential to be generalized to single-organ registration as long as the point clouds have different semantic labels (such as using vessels inside the liver as additional labels). Finally, the homogeneous linear elastic energy is a simplified model, and future work can explore the use of different biomechanical parameters for different control points based on the type of tissue.
@@ 5 Conclusion
We present a novel non-rigid registration method that uses semantic label-based point matching to improve the quality of the matched point pairs. We utilize a novel control point-based deformation representation to enable explicit biomechanics-based regularization in point cloud registration. Our method outperforms state-of-the-art point matching-based registration methods on three different datasets, showing that it is generalizable to different tasks and challenges.

### v3

@@ 3.6 Ablation Studies
W/o semantic matching + RegReg $2.42^{*}$ $90^{*} 0.0$0740.0074
Semantic w/o RegReg $\mathbf{1.15}^{*}$ $\mathbf{0.66}^{*}$ $0.3735^{*}$
Semantic + $Reg_{grad}$ $\mathbf{1.15}^{*}$ $\mathbf{0.66}^{*}$ $0.0137^{*}$
Semantic + $Reg_{mag}$ $\mathbf{1.15}^{*}$ $\mathbf{0.66}^{*}$ $0.0259^{*}$
Semantic + $Reg_{els}$ $\underline{1.17}$ $\underline{0.67}$ $0.0074$
Semantic + $Reg_{grad}+Reg_{mag}$ $\mathbf{1.15}^{*}$ $\mathbf{0.66}^{*}$ $0.0134^{*}$
Semantic + $Reg_{grad}+Reg_{els}$ $\underline{1.17}$ $\underline{0.67}$ $\underline{0.0073}$
Semantic + $Reg_{mag}+Reg_{els}$ $\mathbf{1.15}$ $\underline{0.67}$ $0.0074$
Semantic + $Reg$ (Ours) $\underline{1.17}$ $\underline{0.67}$ $\mathbf{0.0072}$
Figure 7: (a) Comparison of SemICP with and without semantic matching. Left: With semantic+RegReg (axial and coronal views); Right: Without semantic+RegReg (axial and coronal views). The red contour shows the MR segmentation boundary, and the colored mask represents the registered CT. (b) Comparison of deformation fields under different regularization terms. Yellow indicates larger deformation amplitude; contours represent the MR organ segmentation. The elastic regularization achieves a smoother change of the amplitude in the deformation field compared with gradient or magnitude regularization. (Top row: axial view; bottom row: coronal view.)
@@ 3.7 Sensitivity on Elasticity Values
To evaluate the sensitivity of our method to the Young’s modulus $E$ and Poission’s ratio $\nu$, we perform SemICP on the AbdominalMRCT dataset with varying elastic values. Based on finite element simulation (Brock et al. (2005)), typical $E$ for soft tissues and abdominal organs ranges from 1kPa to 100 kPa, except only the stomach which reaches 500kPa, and $\nu$ typically ranges from 0.400 to 0.499. In $E$ sensitivity analysis, we fixed $\nu=0.499$ and varied $E$ across {1kPa, 10kPa, 100kPa, 500kPa, 1MPa}. To also understand the influence of $\nu$, we fixed $E$ at 1kPa and varied $\nu$ in {0.350, 0.400, 0.450, 0.499}.
The sensitivity to Young’s modulus $E$ is presented in Table 7. As $E$ increases, the surface matching errors HD95 and MSD increase slightly, while the spatial distortion SDLogJ decreases. However, the absolute differences across metrics remain small, particularly between 1kPa and 100kPa. The results indicate that as the tissue stiffness increases, the estimated deformation becomes more regularized, but its ability to accurately fit the boundary conditions slightly diminishes. However, in the soft tissues and soft organs range, the differences in performance are small. The sensitivity to Poisson’s Ratio $\nu$ is shown in Table 8. The surface matching has a slight improvement when the $\nu$ decreases, while the SDLogJ increases. It indicates that assuming greater tissue compressibility weakens the linear energy regularization, allowing larger but less regularized deformations to fit the boundary conditions. Nonetheless, the differences in metrics remain small.
@@
HD95 $\downarrow$ 1.17 1.20* 1.23* 1.49 1.85*
MSD $\downarrow$ 0.67 0.68* 0.71* 0.81 0.91*
SDLogJ $\downarrow$ 0.0074 0.0053* 0.0034* 0.0024* 0.0022*
@@ 3.8 Sensitivity on Misalignment and Visualizable Ratio
To assess the sensitivity of our method and its robustness with respect to initial misalignments, we simulated partial observations and applied different rigid misalignments to the AbdominalMRCT dataset. We cropped the original point clouds to partially visible point clouds with 5%, 10%, 25% and 50% of points visible, and introduced additional Gaussian noise. We also added different levels of misalignment, with Euler rotation angles within $\pm 5^{\circ}$, $\pm 15^{\circ}$, and $\pm 30^{\circ}$. For each experiment, we generated 10 different simulated partial point clouds for each patient, resulting in a total of 80 synthetic point cloud pairs. We compare our SemICP-rigid method with ICP and its variants to show how the semantic matching improves ICP robustness. FR-ICP is excluded from comparison due to its higher HD95 in Table 2, indicating less favorable performance. Therefore, we limited the comparison to ICP and GO-ICP. The point clouds are not downsampled before running ICP, GO-ICP or semICP-rigid in this experiment.
As shown in Table 9, SemICP-rigid achieves lower HD95, MSD, and TRE compared to ICP and GO-ICP under challenging visibility conditions (5%, 10%, and 25%). When evaluating sensitivity to initial misalignment, both ICP and GO-ICP exhibit significantly increased registration error at rotation angles of $\pm 30^{\circ}$, indicating sensitivity to initialization. In contrast, SemICP-rigid demonstrates improved robustness, with smaller error increases under the same conditions. These results highlight that incorporating semantic labels enhances the accuracy and stability of the rigid registration step, even under limited visibility and challenging initial alignment.
@@ 4 Discussion
Our work introduces a novel biomechanical energy-based regularization into point-matching-based point cloud registration. Existing methods (NR-ICP, AMM-NRR and Fast-RNRR) regularize the deformation with a local affine regularization to constrain deformation, which fails to capture the underlying biomechanical properties of medical point clouds. Moreover, these methods focus solely on surface matching, neglecting the fact that medical point clouds often undergo deformation as cohesive volumetric structures. This limits their ability to assess and preserve the underlying deformation topology. Our elastic energy-based regularization is physically motivated by the biomechanical properties of anatomical structures, and the ablation studies confirm that the inclusion of the proposed elastic energy regularization leads to significant improvements in the plausibility and quality of the estimated deformation fields, as evidenced by reduced SDLogJ, with little degradation in the surface matching accuracy.
We do not apply the actual elasticity values in the regularization, but the sensitivity studies show that the differences in HD95, MSD and SDLogJ remain small with various $E$ and $\nu$. The non-rigid cost function includes two parts: the first part is boundary conditions that drive the alignment of the organ surfaces, and the second part is the regularization that includes elastic energy. They work together to reach the equilibrium between accuracy and smoothness. Our sensitivity experiments on $E$ show that the registration is more likely to be driven by boundary conditions than by the regularization, particularly when $E$ is within the range of 1kPa to 100kPa, including the soft organs of interest like liver and kidney (Kruse et al. (2000)) and brain (Murphy et al. (2019)). Similar results are observed in finite element-based registration. Chi et al. (2006) found that a 30% material uncertainty results in registration errors within 1.3mm for thin and hollow structures, and though the maximum error for prostate can reach up to 4.5mm, most of the subvolumes have a much smaller registration error. Our sensitivity experiments on $\nu$ indicate that $\nu$ has a small influence on the registration results, which is similar to the finding in the finite element-based registration (Brock et al. (2005)) that varying $\nu$ only leads to small changes in accuracy. In the future, inhomogeneous $E$ and $\nu$ can be employed based on the labels or information from elastography (Nir et al. (2013)). More complicated non-linear effects might be integrated, but it requires the consideration of the trade-off between algorithm complexity and accuracy. A research in finite element-based lung registration shows that the registration error is not significantly affected by the element type (elastic and hyperelastic) or linearity of analysis, suggesting the use of linear finite element analysis for simplicity (Al-Mayah et al. (2011)).
@@
The deep learning-based methods, DCPv1 and DCPv2, increase the surface alignment error in the datasets. The results demonstrate that models trained on daily object data can not be utilized in the medical point clouds in an out-of-the-box manner, which is consistent with findings in previous works (Weber et al. (2024); Zhang et al. (2024)). It is possible that the ModelNet40 data of daily objects typically include clear feature points, such as corners and lines, to assist the alignment, while the medical surfaces typically exhibit smoother local curvature and can be partially visible. As a results, models trained on regular objects such as ModelNet40 can not learn the features that generalize well to medical data. Our non-learning, semantic-aware registration framework circumvents this limitation by directly leveraging anatomical labels and biomechanical constraints without relying on large-scale, object-domain pretraining.
Our work requires semantic labels in the point cloud, and this paper focuses on registration and does not address the segmentation process in detail. We present a proof-of-concept study of integrating a deep learning model to segment the volumes from different modalities and to perform the registration on the imperfectly segmented point clouds. Our experiments show that our method can register the less ideal point clouds, supporting its feasibility for integration into US–based image guidance systems, where segmentation and registration must operate on noisy or incomplete data. Moreover, we highlight that our control-grid-based deformation enables warping the ground truth point cloud labeled by human experts. It is almost impossible to transfer the registration results from the segmented point cloud pairs to ground truth point cloud pairs using other non-rigid point-matching-based registration, since it is difficult to interpolate the motion of points from the sparse surface point correspondences. Especially when the segmentation point clouds have mis-segmentation, the motion in the unobserved region can not be approximated easily. Our method, with an explicit deformation field, can estimate the movement of unseen points within the deformation field through interpolating their motions based on control grids. The method makes it possible to estimate the motion of new or unobserved regions, even when segmentation inaccuracies are present. This proof of concept demonstrates a possible pathway toward integrating learning-based segmentation and non-rigid point cloud registration into US-guided intervention workflows.
This method is designed for multi-organ registration, but it has the potential to be generalized to single-organ registration as long as the point clouds have different semantic labels (such as using vessels inside the liver as additional labels). Additionally, the homogeneous linear elastic energy is a simplified model, and future work can explore the use of different biomechanical parameters for different control points based on the type of tissue.
@@ 5 Conclusion
We present a novel non-rigid point cloud registration method that extends semantic label-based point matching to improve the quality of the matched point pairs in medical point clouds. We utilize a novel control point-based deformation representation to enable explicit biomechanics-based regularization in point cloud registration. Our method outperforms state-of-the-art point-matching-based registration methods on three different datasets, showing that it is generalizable to different tasks and challenges. Through a proof-of-concept study, we demonstrate that integrating deep learning segmentation with our registration pipeline enables effective alignment of imperfectly segmented US and MR point clouds. These results highlight the potential of our framework to be used in CAI systems.
CRediT authorship contribution statement
Wanwen Chen: Conceptualization, Methodology, Software, Visualization, Investigation, Data curation, Formal analysis, Writing - original draft; Qi Zeng: Software, Formal analysis, Data curation, Investigation, Writing - review and editing; Carson Studders: Software, Visualization, Data curation, Writing - review and editing; Jamie J.Y. Kwon: Investigation, Data curation; Emily H.T. Pang: Investigation, Data curation, Supervision; Eitan Prisman: Conceptualization, Investigation, Data Curation, Funding acquisition, Supervision, Resources; Septimiu E. Salcudean: Conceptualization, Methodology, Funding acquisition, Supervision, Resources, Writing - review and editing.
Declaration of competing interest
The authors have no competing interests to declare that are relevant to the content of this article.
Acknowledgment
The work is supported by an NSERC Discovery Grant and the Charles Laszlo Chair in Biomedical Engineering held by Dr. Salcudean, and by the VCHRI Innovation and Translational Research Award, the University of British Columbia Department of Surgery Seed Grant and CIHR Project Grant held by Dr. Prisman.
Data statement
Data from the TORS study and the liver study are not openly available due to the ethical restrictions, that the data contain sensitive patient information. Learn2reg challenge datasets are available through accessing the challenge website https://learn2reg.grand-challenge.org/Datasets/. Learn2reg AbdomenMRCT dataset is openly available in https://cloud.imi.uni-luebeck.de/s/32WaSRaTnFk2JeT. Learn2reg hippocampus dataset is openly available in https://cloud.imi.uni-luebeck.de/s/gPN6bzJ4xf7HzG8.
Declaration of generative AI and AI-assisted technologies in the manuscript preparation process
Statement: During the preparation of this work, the author(s) used ChatGPT for grammar correction and language editing. After using this tool/service, the author(s) reviewed and edited the content and take(s) full responsibility for the content of the published article.